find x
1350=0.25x+1750-.035x
show work

Answer by mananth(16949)

1350=0.25x+1750-.035x
multiply by 1000 to remove the decimals
1350000=250x+1750000-35x
add -1750000 on both sides
1350000-1750000= 215x
-40000=215x
/215
-40000/215 =x
x=1860.46
...
m.ananth@hotmail.ca